OpenStudy (anonymous):

16L of antifreeze is 30% antifreeze how much mixture should be drained and replaces with pure antifreeze so final mix is 50% anitfreeze

OpenStudy (anonymous):

Let x be amount removed to make way for pure, then 30(16-x)/100 + 100x/100 =50*16/100 -> 480 + 70x = 800 -> x = 320/70 = 32/7 litres to remove

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

are you familiar with the greek letter "mu"\[\mu\]? it look like a population problem but its got that thing in it I=I\[I=I_{0}e ^{-mux} \mu=1.4\] percentage of light intensity at 1m,3m,5m,10m, 50m?

OpenStudy (anonymous):

I don't really know, scouting about a bit seems to indicate it is used in descriptions of light intensity as u suggest. Eg 86 mu mol m^-2 s^-1 which I would be inclined to read as micro mol but that's just a guess I'm afraid.